# Adaptive Partition Factor Analysis

### v2

@@ 2.2 Model properties
 $\>1=\|P_{{s^{\prime}},h}\|^{2}=\sum_{l\in A_{s^{\prime}}\cap A_{s}}P_{lh}^{2}+\sum_{l\in A_{s^{\prime}}\setminus A_{s}}P_{lh}^{2}.$ (10)
Equations (8), (9), and (10) enforce that all $P$, $P_{s}$, and $P_{s^{\prime}}$ are orthogonal matrices. These set of equations take the form: $a+b+c+d=1,\>a+b=1,\>a+c=1$ with $a,b,c,d\geq 0$. This system admits only the solution $a=1$ and $b=c=d=0$. This observation directly constrains the entries of the $h$-th column of $P$ pertaining to the last three summations of (8) to be zero. Consequently, the only actual rotations, i.e. those induced by non-diagonal $P_{s}$, are those corresponding to factors shared between groups $s$ and $s^{\prime}$, i.e. the $h$-th column of $P$ must vanish outside $A_{s}\cap A_{s^{\prime}}$.
Assumption 2 induces further constraints on $P$. Specifically, for any pair $(h,h^{\prime})$ jointly in groups $s$ and $s^{\prime}$, there exists a third group $s^{\prime\prime}$ such that either $h\in A_{s^{\prime\prime}}$ but $h^{\prime}\notin A_{s^{\prime\prime}}$, or $h^{\prime}\in A_{s^{\prime\prime}}$ but $h\notin A_{s^{\prime\prime}}$. Applying the same norm argument to $s$ and $s^{\prime\prime}$ (or equivalently to $s^{\prime}$ and $s^{\prime\prime}$) forces the $h$th column of $P$ to vanish outside $A_{s}\cap A_{s^{\prime\prime}}$. Crucially, since $A_{s}\cap A_{s^{\prime\prime}}$ excludes $h^{\prime}$ (or $h$), we conclude $P_{hh^{\prime}}=P_{h^{\prime}h}=0$ and $P_{hh}=P_{h^{\prime}h^{\prime}}=\pm 1$. The argument is valid for all $h$, leading to $P$ being a a diagonal matrix with entries ±$\pm 1$. ∎
Clearly, the shared variability induced by $\Lambda$ remains susceptible to well-known identifiability issues. However, these challenges can be effectively addressed using state-of-the-art methods, such as those proposed by Poworoznek et al. (2021) and Papastamoulis & Ntzoufras (2022).
@@ 3 Simulation study
We assess the performance of APAFA in a simulation experiment and its relative merits with respect to the approach of Grabski et al. (2023) (TETRIS).
We consider two main configurations concerning the data dimensionality: tall data where $n=60$ and $p=10$ and large data where $n=45$ and $p=60$ and for each of them, simulate data from different scenarios. Scenario A represents the case in which the original MSFA model is correctly specified. Namely, we have $S=3$ groups with sample sizes $n_{1}=n_{2}=n_{3}=n/3$ and each of them has a single different study-specific factor. The total number of shared factor is $k=3$. The groups labels match exactly the three groups. To test our model’s ability to detect the underlying structure, we also fit the model without providing the labels—meaning the model is unaware of the existence of three distinct groups. We refer to this situation as Scenario $\text{A}^{*}$. In Scenario B there are $S=3$ studies that are perfectly homogeneous, i.e. $k_{s}=0$ for each $s=\{1,2,3\}$ or equivalently $\Phi$ has all entries equal to zero. Notably, we fit each competing model providing the information that $S=3$ groups are present. Scenario C presents a complex situation in which there exist $S=3$ studies with partially shared latent factors. Specifically, study 1 and 2 present their own factors ($\varphi_{1},\varphi_{2})$ as in Scenario $A$, while a third factor is present for units of the second and the third study. Scenario D addresses a situation in which for each of the three studies a subset of units has no specific factors. The left panels of Figure 3 provide a representation of the true specific latent factors sparsity pattern. Further details on the data-generating process are reported in the Supplementary Materials.
Each combination of scenarios and dimensions is replicated independently $R=10$ times. For each of them, all the competing methods have been fit. For the scenario where $n>p$, the computation time for a single replicate ranged from 3 to 4 minutes, whereas for the case $p>n$, it spanned from 10 to 13 minutes on a laptop CPU with a clock speed of 1.00 GHz. More details on the computing time are available in the Supplementary Materials 1. Posterior estimates for all methods are obtained by running MCMC algorithms for each of the $R$ replicates for 10,000 iterations, discarding the first 8,000, and then keeping the last 2,000 values. An implementation of the Gibbs sampler reported in the Supplementary Materials is available at https://github.com/elenabortolato/APAFA. For TETRIS we used the implementation available at https://github.com/igrabski/tetris.
@@
Tables 1 and 2 report the results. Table 1 shows that both methods are comparable in terms of posterior concentration around the true number of factors with a slightly better performance by APAFA. Notably, TETRIS has not been fitted in scenario A, since no labels are assigned to units, and in scenario D since the method is not suited to find subgroups of units with different factor within a given group. In terms of estimation of the shared variance part, TETRIS has a slightly better or comparable performance in Scenario A when it is correctly specified but has an overall worse performance in all the other cases. Qualitatively similar results are observed in the $n<p$ case, reported in Table 2. The performance of estimation of the shared variance, often the interest of the analysis, is reported in Figure 4. In Scenario B, characterized by the absence of specific factors, TETRIS successfully recognizes the shared structure. However, it faces challenges in distinguishing the contributions to the variance components when group-specific characteristics are present.
@@
Figure 5: ROC curves (left) and distribution of the AUC (right) computed over 10 experiments for the posterior probability of assignment of the specific latent factors to units in for each scenario (from top to bottom: A, A, C, D), with $n>p$.
To assess the sparsity of the specific factors and the induced adaptive partitioning of units into subgroups that do not deterministically align with the grouping variable $x$, we compared the true zero-one pattern used to simulate the synthetic data reported in Figure 3 with the estimated $\psi_{ih}(x_{i})$. To address the issue of column permutations, for sake of this simulation evaluation when the true value of $\Phi$ is known, we post-processed the order of the estimates of $\Phi$. Details are reported in the Appendix. Figure 5 displays the ROC curves along with the corresponding area under the curve (AUC) for each replicated dataset. The results indicate a strong ability of APAFA to detect the true partition induced by the sparsity pattern of the specific factors in $\Phi$ in each scenario.
@@ 4.1 Bird species occurence dataset
We first examine the co-occurrence patterns of $p=50$ bird species in Finland analyzing a data set collected over nine years (2006-2014) across $S=200$ locations (Lindström et al., 2015). Notably, part of this datased has been already analyzed in Schiavon et al. (2022). The average number of sightings per location in the entire period is of about 5, for a total of $n=914$ sites examined during the years. Shared factors may depend on ambient characteristics as temperature, latitude, habitat type, and proximity to the ocean. To illustrate the performance of the proposed model in this context we avoid a fine model specification including these covariates and rather see if a the estimated latent factors are able to reconstruct some of these information.
@@
We set the hyperparameters governing a priori the number of active common and specific factors to $\alpha^{\eta}=15$ and $\alpha^{\varphi}=15$ respectively. We run the Gibbs sampler for 10,000 iterations and use the last 5,000 to obtain posterior estimates.
A posteriori, the estimated number of active factors are $\hat{d}=5$ and $\hat{k}=5$. Among all the possible aspects, it is interesting to focus on the interpretation of the main specific factors $\varphi_{i1}$. Notably, this factor takes non-zero values at a subset of locations, effectively grouping observations across multiple sites. This behavior aligns with the possibility of a partially-shared latent factor similarly to Grabski et al. (2023). Interestingly, a post-hoc analysis of these locations, using a categorical covariate that describes habitat type (not used in the model fitting), revealed that these locations are predominantly urban. Consistently with this, it is clear that the proposed method is able to find subject-specific factors associated with unobserved grouping variables. See the first panel of Figure 6 where the specific factor matrix $\Phi$ is reported with the columns ordered by habitat type. Notably, this finding aligns with recent ornithological literature examining the impact of urbanization on bird interaction (Pena et al., 2023). Similar qualitative insights can be appreciated in the second panel of the same figure where the columns are ordered with the associated latitude value. Notably, both the first and fourth factors are strongly associated with latitude.
@@ 4.2 Immune dataset
As a second example, we analyze a publicly accessible dataset which contains transcriptomic data for $p=63$ genes linked to immune system function in female oncology patients with ovarian cancer. This dataset can be accessed through the curatedOvarianData package in Bioconductor Ganzfried et al. (2013). Notably, this same dataset was utilized in the seminal MSFA paper by De Vito et al. (2019). This dataset comprises four studies: GSE9891 and GSE20565, which utilize the same microarray platform for data acquisition, as well as TCGA and GSE26712, with respective sample sizes of 285, 140, 578, and 195 units ($n=1198$ observations in total). We fit model (2) with hyperparameters governing a priori the number of active common and specific factors to $\alpha^{\eta}=1$ and $\alpha^{\varphi}=4$, respectively. The chain was initialized from $d=k=12$ active factors. The shape and rate parameters chosen for the Inverse Gamma prior for the elements of $\Lambda,\Gamma$ and diagonal elements of $\Sigma$ were all equal to 2. We run the Gibbs sampler for 10,000 iterations and use the last 5,000 to obtain posterior estimates.
Posterior analysis identified 5 shared factors and 8 specific factors. Notably, the number of shared factors is consistent with the findings of De Vito et al. (2019) using the BIC criterion. As in the previous section, we compared the resulting matrix of activated specific factors to covariate data not included in the initial analysis, which are also available in the curatedOvarianData package. This comparison provides additional validation and insight into the biological relevance of the factors identified. Figure 7 shows that the grouping associated with the studies is captured within the structure of $\Phi$. Specifically, factors $\varphi_{6}$ and $\varphi_{7}$ are clearly related to the group structure provided by $x$, in particular with the TCGA and Gaus experiments, while individuals from the GSE9891 and GSE20565 studies, both using the same microarray platform for data processing, do not present specific variance adjustments. The remaining factors are linked to a small number of units each, resulting in numerous configurations, far exceeding the initially presumed four. Upon examining external covariates not included in the analysis, we found that $\varphi_{1}$ is associated with 6 units that are distinct from the rest, as they exhibit the clear cell histological subtype. Factor 3 distinguishes the gene expression of 7 patients with mucinous carcinoma from other types of carcinoma and 8 patients with tumors in the fallopian tubes (ft) rather than in the ovary. Factors 4 and 8 discriminate patients with a tumor classified as other (sarcomatoid, adenocarcinoma, dysgerminoma) together with some units presenting clearcell and endometrial tumor (endo, 26 units). See the boxplots in Figure 7 for details.
@@
Figure 8 illustrates the contributions to the correlation matrix given by $\Gamma_{h}\Gamma_{h}^{\top}$ for $h=1,\ldots,8$. The comparison with Tetris, eliciting a prior coherent with that used in APAFA, i.e. 5 factors in total, with 4 specific, revealed instead 46 active factors, with 90% credible interval 29-63. Among these, 26 are common and 20 specific.
Finally, to evaluate the model’s goodness of fit, we run the Gibbs sampler for APAFA and TETRIS a second time, excluding 30 observations and imputing those missing values using the associated posterior predictive distributions. The mean squared errors were 0.0451 and 0.2833 for APAFA and TETRIS, respectively. Further details are reported in the Supplementary Materials.
@@ 5 Discussion
Motivated by the strict grouping imposed by MSFA approaches, we introduced a factor analytic model where the partition induced by specific latent factors is influenced by external information but is not strictly determined. While the primary interest has been in using a grouping categorical variable, consistently with the study structure of MSFA, the proposed APAFA is much more general and offers the possibility to include also other categorical or continuous covariates. These act in promoting sparsity in the latent factors, as done by Schiavon et al. (2024) thereby providing a natural way to aid in the modelization of the covariance structure and connecting with modelling the heterogeneity in factor regression (Avalos-Pacheco et al., 2022).
An important aspect of our contribution is that it is not covariate dependence per se that distinguishes specific from shared factors, but rather the sparsity caused by this dependence. Accordingly, there is no fundamental barrier to shared factors depending common covariates, as long as such dependence does not lead to sparse latent factors. In principle, one could consider an alternative model in which shared factors vary with the covariates in a non-sparse manner. While exploring such an extension is conceptually feasible, it lies beyond the scope of this paper, which focuses on proposing an innovative alternative to the standard MSFA framework. We leave these broader generalizations to future work.
The proposed solutions proved to be beneficial across many simulation settings, even in cases with group misspecification and complex heterogeneity both between and within the fixed groups. Overlapping group structures are nonetheless detectable and may differ in terms of flexibility and granularity. For example, in Section 4.1, study units are grouped together in scenarios where the number of studies closely matches the number of units. Conversely, in the example of Section 4.2 the model enables detecting higher granularity with respect to that specified a priori, identifying clusters of a dozen units out of a thousand.
@@
Details on simulation study
Details on the data generating mechanism, prior hyperparameters, initialization strategy used, sensitivity analysis and convergence assessments in the simulation experiments and real data analyses of Section 3 (pdf file).
Details on theory and computation
A detailed description of the Gibbs sampler algorithm derived to perform posterior computation is provided. A short description of the reordering criterion applied to the estimated factors to evaluate the quality of the partitioning into subgroups in the experiments of Section 3. Proofs complementing Section 4 (pdf file).

### v3

@@ 2.3 Identification properties
 $1=\|P_{{s^{\prime}},h}\|^{2}=\sum_{l\in A_{s^{\prime}}\cap A_{s}}P_{lh}^{2}+\sum_{l\in A_{s^{\prime}}\setminus A_{s}}P_{lh}^{2}.$ (10)
Equations (8), (9), and (10) enforce that all $P$, $P_{s}$, and $P_{s^{\prime}}$ are orthogonal matrices. These set of equations take the form: a+b+c+d=1,a+b=1,a+c=1a+b+c+d=1,\>a+b=1,\>a+c=1 with $a,b,c,d\geq 0$. This system admits only the solution $a=1$ and $b=c=d=0$. This observation directly constrains the entries of the $h$-th column of $P$ pertaining to the last three summations of (8) to be zero. Consequently, the only actual rotations, i.e. those induced by non-diagonal $P_{s}$, are those corresponding to factors shared between groups $s$ and $s^{\prime}$, i.e. the $h$-th column of $P$ must vanish outside $A_{s}\cap A_{s^{\prime}}$.
Assumption 2 induces further constraints on $P$. Specifically, for any pair $(h,h^{\prime})$ jointly in groups $s$ and $s^{\prime}$, there exists a third group $s^{\prime\prime}$ such that either $h\in A_{s^{\prime\prime}}$ but $h^{\prime}\notin A_{s^{\prime\prime}}$, or $h^{\prime}\in A_{s^{\prime\prime}}$ but $h\notin A_{s^{\prime\prime}}$. Applying the same norm argument to $s$ and $s^{\prime\prime}$ (or equivalently to $s^{\prime}$ and $s^{\prime\prime}$) forces the $h$th column of $P$ to vanish outside $A_{s}\cap A_{s^{\prime\prime}}$. Crucially, since $A_{s}\cap A_{s^{\prime\prime}}$ excludes $h^{\prime}$ (or $h$), we conclude $P_{hh^{\prime}}=P_{h^{\prime}h}=0$ and $P_{hh}=P_{h^{\prime}h^{\prime}}=\pm 1$. The argument is valid for all $h$, leading to $P$ being a a diagonal matrix with entries $\pm 1$. ∎
Clearly, the shared variability induced by $\Lambda$ remains susceptible to well-known identifiability issues. These challenges can be effectively addressed using state-of-the-art methods for post-processing the MCMC output such as those proposed by Poworoznek et al. (2025) and Papastamoulis & Ntzoufras (2022) or by imposing additional structural constraints on the loading matrix, as in Frühwirth-Schnatter et al. (2023, 2025). Interestingly, the constraints introduced in the latter works are formulated in terms of sparsity in the loading matrix. Although our identifiability conditions are imposed on the factors, whereas Frühwirth-Schnatter et al. (2023) focus on the factor loadings, both approaches exploit the concept of sparsity, highlighting its central role in achieving identifiability in factor models.
@@ 3 Simulation study
We assess the performance of APAFA in a simulation experiment and its relative merits with respect to the approach of Grabski et al. (2023) (TETRIS). We consider two main configurations concerning the data dimensionality: tall data where $n=60$ and $p=10$ and large data where $n=45$ and $p=60$ and for each of them, simulate data from different scenarios. Scenario A represents the case in which the original MSFA model is correctly specified. Namely, we have $S=3$ groups with sample sizes $n_{1}=n_{2}=n_{3}=n/3$ and each of them has a single different specific factor. The total number of shared factor is $k=3$. The groups labels match exactly the three groups. To test our model’s ability to detect the underlying structure, we also fit the model without providing the labels—meaning the model is unaware of the existence of three distinct groups. We refer to this situation as Scenario $\text{A}^{*}$. In Scenario B there are $S=3$ studies that are perfectly homogeneous, i.e. $k_{s}=0$ for each $s=\{1,2,3\}$ or equivalently $\Phi$ has all entries equal to zero. Notably, we fit each competing model providing the information that $S=3$ groups are present. Scenario C presents a complex situation in which there exist $S=3$ studies with partially shared latent factors. Specifically, study 1 and 2 present their own factors ($\varphi_{1},\varphi_{2})$ as in Scenario A, while a third factor is present for units of the second and the third study. Scenario D addresses a situation in which for each of the three studies a subset of units has no specific factors. The left panels of Figure 3 provide a representation of the true specific latent factors sparsity pattern. Further details on the data-generating process are reported in the Supplementary Material.
Each combination of scenarios and dimensions is replicated independently $R=10$ times. For each of them, all the competing methods have been fit. For the scenario where $n>p$, the computation time for a single replicate ranged from 3 to 4 minutes, whereas for the case $p>n$, it spanned from 10 to 13 minutes on a laptop CPU with a clock speed of 1.00 GHz. More details on the computing time are available in the Supplementary Material. Posterior estimates for all methods are obtained by running MCMC algorithms for each of the $R$ replicates for 10,000 iterations, discarding the first 8,000, and then keeping the last 2,000 values. An implementation of the Gibbs sampler reported in the Supplementary Material is available at https://github.com/elenabortolato/APAFA. For TETRIS we used the implementation available at https://github.com/igrabski/tetris.
@@
Tables 1 and 2 report the results. Table 1 shows that both methods are comparable in terms of posterior concentration around the true number of factors with a slightly better performance by APAFA. Scenario A is the most favorable for TETRIS, as the model is correctly specified. Indeed, although TETRIS struggles to disentangle the specific and shared components, it still provides overall good estimates of the covariance matrices with slightly better or comparable performance to APAFA. Considering the standard errors, such superiority is not striking. TETRIS has not been fitted in scenario A, since no labels are assigned to units, and in scenario D, since the method is not suited to find subgroups of units with different factors within a given group.
Qualitatively similar results are observed in the $n<p$ case, reported in Table 2. The performance of estimation of the shared variance, often the interest of the analysis, is reported in Figure 4. In Scenario B, characterized by the absence of specific factors, TETRIS successfully recognizes the shared structure. However, it faces challenges in distinguishing the contributions to the variance components when group-specific characteristics are present.
@@
Figure 5: ROC curves (left) and distribution of the AUC (right) obtained from posterior probabilities under the APAFA model over 10 independent replicated datasets, illustrating the accuracy of factor-to-unit assignments across scenarios (from top to bottom: A, A, C, D), with $n>p$.
To assess the sparsity of the specific factors and the induced adaptive partitioning of units into subgroups that do not deterministically align with the grouping variable $x$, we compared the true zero/one pattern used to simulate the synthetic data reported in Figure 3 with the estimated $\psi_{ih}(x_{i})$. To address the issue of column permutations, for the sake of this simulation evaluation, when the true value of $\Phi$ is known, we post-processed the order of the estimates of $\Phi$. Details are reported in the Appendix. Figure 5 displays the ROC curves along with the corresponding area under the curve (AUC) for each replicated dataset. The results indicate a strong ability of APAFA to detect the true partition induced by the sparsity pattern of the specific factors in $\Phi$ in each scenario.
Additional assessments, including a sensitivity analysis and a simulation with nearly identical study-specific factors, further confirm APAFA’s empirical performance (see Supplementary Materials for details).
@@ 4.1 Bird species occurence dataset
We first examine the co-occurrence patterns of $p=50$ bird species in Finland analyzing a data set collected over nine years (2006-2014) across $S=200$ locations (Lindström et al., 2015). The average number of sightings per location during the entire period is about 5, for a total of $n=914$ sites examined over the years. Shared factors may depend on ambient characteristics such as temperature, latitude, habitat type, and proximity to the ocean. To illustrate the performance of the proposed model in this context, we avoid fine model specification including these covariates, and rather examine if the estimated latent factors are able to reconstruct some of this information.
@@
We set the hyperparameters governing a priori the number of active common and specific factors to $\alpha^{\eta}=15$ and $\alpha^{\varphi}=15$ respectively. We run the Gibbs sampler for 10,000 iterations and use the last 5,000 to obtain posterior estimates. Five active factors were identified in both shared and specific components. Among the various aspects, we focus on interpreting the main specific factors, $\varphi_{i1}$. This factor takes non-zero values at a subset of locations, effectively grouping observations across multiple sites. This behavior aligns with the possibility of a partially-shared latent factor, similarly to Grabski et al. (2023). A post-hoc analysis of these locations, using a categorical covariate that describes habitat type (not used in the model fitting), revealed that these locations are predominantly urban. Consistently with this, it is clear that the proposed method is able to find subject-specific factors associated with unobserved grouping variables. See the first panel of Figure 6 where the specific factor matrix $\Phi$ is reported with the columns ordered by habitat type. Notably, this finding aligns with recent ornithological literature examining the impact of urbanization on bird interaction (Pena et al., 2023). Similar qualitative insights can be appreciated in the second panel of the same figure, where the columns are ordered by the associated latitude values. Interestingly, both the first and fourth factors are strongly associated with latitude. Uncertainty quantification for these figures is provided in the Supplementary Material.
@@ 4.2 Immune dataset
As a second example, we analyze a publicly accessible dataset that contains transcriptomic data for $p=63$ genes linked to immune system function in female oncology patients with ovarian cancer. This dataset can be accessed through the curatedOvarianData package in Bioconductor Ganzfried et al. (2013). The same dataset was also utilized in the seminal MSFA paper by De Vito et al. (2019). This dataset comprises four studies: GSE9891 and GSE20565, which utilize the same microarray platform for data acquisition, as well as TCGA and GSE26712, with respective sample sizes of 285, 140, 578, and 195 units ($n=1198$ observations in total). We fit model (2) with hyperparameters governing a priori the number of active common and specific factors to $\alpha^{\eta}=1$ and $\alpha^{\varphi}=4$, respectively. The chain was initialized from $d=k=12$ active factors. The shape and rate parameters chosen for the Inverse Gamma prior for the elements of $\Lambda,\Gamma$ and the diagonal elements of $\Sigma$ were all equal to 2. We run the Gibbs sampler for 10,000 iterations and use the last 5,000 to obtain posterior estimates.
Posterior analysis identified 5 shared factors and 8 specific factors. Notably, the number of shared factors is consistent with the findings of De Vito et al. (2019) using the BIC criterion. As in the previous section, we compared the resulting matrix of activated specific factors to covariate data not included in the initial analysis, which are also available in the curatedOvarianData package. This comparison provides additional validation and insight into the biological relevance of the factors identified. Figure 7 shows that the grouping associated with the studies is captured within the structure of $\Phi$. Specifically, factors $\varphi_{6}$ and $\varphi_{7}$ are clearly related to the group structure provided by $x$, in particular with the TCGA and Gaus experiments, while individuals from the GSE9891 and GSE20565 studies, both using the same microarray platform for data processing, do not present specific variance adjustments. The remaining factors are linked to a small number of units each, resulting in numerous configurations that far exceed the initially presumed four. Upon examining external covariates not included in the analysis, we found that $\varphi_{1}$ is associated with 6 units that are distinct from the rest, as they exhibit the clear cell histological subtype. Factor 3 distinguishes the gene expression of 7 patients with mucinous carcinoma from other types of carcinoma and 8 patients with tumors in the fallopian tubes (ft) rather than in the ovary. Factors 4 and 8 discriminate patients with tumors classified as other (sarcomatoid, adenocarcinoma, dysgerminoma), along with some units presenting clear cell and endometrial tumors (endo, 26 units). See the boxplots in Figure 7 for details.
@@
Figure 8 illustrates the contributions to the correlation matrix given by $\Gamma_{h}\Gamma_{h}^{\top}$ for $h=1,\ldots,8$. The comparison with Tetris, eliciting a prior coherent with that used in APAFA, i.e., 5 factors in total, with 4 specific, revealed instead 46 active factors, with a 90% credible interval of 29-63. Among these, 26 are common, and 20 are specific. To evaluate the model’s goodness of fit, we run the Gibbs sampler for APAFA and TETRIS a second time, excluding 30 observations and imputing the missing values using the associated posterior predictive distributions. The mean squared errors were 0.0451 and 0.2833 for APAFA and TETRIS, respectively. Further details are reported in the Supplementary Material.
@@ 5 Discussion
Motivated by the strict grouping imposed by MSFA approaches, we introduced a factor analytic model where the partition induced by specific latent factors is influenced by external information but is not strictly determined. While the primary interest has been in using a grouping categorical variable, consistent with the study structure of MSFA, the proposed APAFA is much more general and offers the possibility of including other categorical or continuous covariates. These act to promote sparsity in the latent factors, as done by Schiavon et al. (2024), thereby providing a natural way to aid in the modeling of the covariance structure and connecting with the modeling of heterogeneity in factor regression (Avalos-Pacheco et al., 2022).
An important aspect of our contribution is that it is not covariate dependence per se that distinguishes specific from shared factors, but rather the sparsity caused by this dependence. Accordingly, there is no fundamental barrier to shared factors depending on common covariates, as long as such dependence does not lead to sparse latent factors. In principle, one could consider an alternative model in which shared factors vary with the covariates in a non-sparse manner. While exploring such an extension is conceptually feasible, it lies beyond the scope of this paper, which focuses on proposing an innovative alternative to the standard MSFA framework. We leave these broader generalizations to future work.
The proposed solutions proved to be beneficial across many simulation settings, even in cases of group misspecification and complex heterogeneity, both between and within the fixed groups. Overlapping group structures are nonetheless detectable and may differ in terms of flexibility and granularity. For example, in Section 4.1, study units are grouped together in scenarios where the number of studies closely matches the number of units. Conversely, in the example of Section 4.2, the model enables detecting higher granularity with respect to that specified a priori, identifying clusters of a dozen units out of a thousand.
@@
The supplementary material provides a detailed description of the Gibbs sampler algorithm used for posterior computation, along with a comprehensive overview of the simulation studies and real-data analyses. Specifically, it includes the data-generating mechanism for simulations, justification for prior hyperparameter selection and sensitivity analyzes, initialization strategies for MCMC chains, further simulation results, convergence and mixing assessment, post-processing steps, and runtime details. (pdf file).
ACKNOWLEDGMENTS
The authors are grateful to Roberta De Vito, Giovanni Parmigiani, Lorenzo Schiavon and Alberto Tonolo for comments on an early version of this paper and to the Associate Editor and the Referees for their constructive comments and suggestions. The authors report there are no competing interests to declare.